A tuning fork of frequency 512Hz is used to tune a piano. The beat frequeny 

observed is 8Hz. What is the frequency of the piano ? How do you ascertain 

the answwer you have chosen is the cdorrect one ?. ( NOTE THAT THERE ARE 

TWO PROBABLE ANSWERS HERE. I WANT TO KNOW HOW YOU WILL GO 

ABOUT ASCERTAINING WHICH IS CORRECT?)


1
Expert's answer
2021-03-23T11:11:27-0400
"f_1=512+8=\ 520\ Hz\\f_2=512-8=\ 504\ Hz\\"

All answers are correct.